OpenStudy (anonymous):

what is the slope of 6,4 and3,2

OpenStudy (anonymous):

\[\frac{4-2}{6-3}=\frac{2}{3} \]

OpenStudy (anonymous):

(Difference in the ys)/(Difference in the xs)

OpenStudy (anonymous):

whats the slope of 1,2.5 and -1,-2.5